8) The number of visitors to the zoo and the entrance fee is shown in the table below (10 p): Entrance fee, Euro Number of visitors per day, persons 0 1,000 700 400 100 4 10 13 Calculate and graphically show the change in consumer surplus if the entrance fee increases from thige (4) to-i-Euro (7). (6 p) How does the price increase affect the revenue of the zoo? (4 p)
Answer #1

Answer:

consumer surplus is the different between the amount a consumer is willing to pay and the actual amount he has to pay for a commodity or service.

consumer surplus = amount consumer is willing to pay - actual amount he has to pay

the willingness to pay for the entrance of the zoo is 13 euro for the consumer and when the entrance fee of the zoo is 4 euro the consumer's surplus is = 13 - 4 = 9 euro. as shown in the below graph the shaded area above the 4 euro till the 13 euro is the surplus of the consumer.

13 lR lo 6 lo o 0 oo yoo 500 600 70o 80o 00 (o oo wumber e visi tors per day, 2/

now if the entrance fee increases from 4 euro to 7 euro the surplus of the consumer decreases as shown in the below graph by the shaded area above 7 euro to till 13 euro

IR lo Consumer surplus 100 Qoo Joo 400 500 600 0 o 800 goo (ooo wumber e visi tos per day.

we can clearly see from both the graphs that, with the increase in the entrance fee of the zoo from 4 euro to 7 euro the consumer surplus has decreased

when the entrance fee is 7 euro the consumer surplus is = 13 - 7 = 6 euro

the consumer surplus has decreased by 3 euro as the consumer surplus was 9 euro when the entrance fee was 4 euro and the consumer surplus become 6 euro when the entrance fee increases to 7 euro.

the increase in price has not changed the revenue of the Zoo because the revenue was 2800 euro per day when the entrance fee was 4 euro as 700 people were visiting the zoo per day and when the prices of entry fee increases to 7 euro than the 400 people were visiting the zoo and the revenue was still 2800 euro so the increase in price did not affected the revenue of the zoo.

revenue = price of entrance fee * number of visitors per day

revenue when price was 4 euro = 4 * 700 = 2800 euro

revenue when price increases to 7 euro = 7* 400 = 2800 euro